Circular No. 8301

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004aq IN NGC 4012 Further to IAUC 8296, M. Armstrong reports his discovery of a sup-

ernova located at α = 11^{h}58^{m}27^{s}.25, δ = +10^{o}01′04′′.5, which is 4′′.6 west and 12′′.8 south of the center of NGC 4012. Available unfiltered-CCD magnitude estimates (by Armstrong, unless noted otherwise): 2002 May 12 UT, [20.0; 2004 Mar. 2.128, 17.5; 7.080, 17.5; 8.848, 17.5 (D. Buczynski, Lancaster, U.K.).

N. Elias-Rosa and S. Benetti, Osservatorio Astronomico di Padova;, V. Stanishev and A. Goobar, Stockholm University; and A. Järvinen, Nordic Optical Telescope (NOT), on behalf of the ‘ERTN’ (cf. IAUC 7987), report that inspection of a spectrum (range 320–910 nm, resolution 1.7 nm), obtained on Mar. 10.01 UT with the NOT (+ ALFOSC) at La Palma, that SN 2004aq is a type-II supernova, ∼ 1 month after explosion. The spectrum is dominated by P-Cyg lines of H, Ca II, and Fe II overimposed on a relatively blue continuum (T_{bb} ∼ 8000 K). The expansion velocities deduced from the Hα and Hβ absorptions are ∼ 7700 and 6500 km/s, respectively.

IRAS 05436–0007 W. D. Vacca and M. C. Cushing, Ames Research Center, NASA; and

J. T. Rayner and T. Simon, University of Hawaii, report on a mediumresolution 0.8–4.9-μm spectrum of the near-infrared counterpart to IRAS 05436–0007 (cf. IAUC 8284), acquired on Mar. 8.25 UT using the NASA Infrared Telescope Facility (+ SpeX facility spectrograph) on Mauna Kea. The data show a broad absorption feature centered at 3.1 μm due to water ice (which has an optical depth of ∼ 0.6) and a deep solid-state absorption at 4.7 μm of frozen CO. Also present are a number of emission lines in the Brackett and Paschen series of H I. The strong emission lines of Paδ and Paγ exhibit prominent P-Cyg components. The CO overtone bands at 2.3 μm appear in emission, as do the Ca II infrared triplet lines at 0.85 μm. Lines of Mg I, Na I, Ca, and Fe are also seen in emission. Several He I lines present in absorption exhibit possible weak P-Cyg structure. The P-Cyg features, plus the appearance of the object in the 2MASS survey at the position of a previously known IRAS and millimeter-wave source (LMZ 12; cf. Lis et al. 1999, Ap.J. 527, 856), suggest that the current activity of the source may arise from the onset of a strong outflow or the recent breakout of a deeply embedded class-I object, rather than from an accretion episode of a young FU-Ori star.

2004 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green